Grade 12th passMechanics

a force F=5t is applied on a lock of mass 10kg kept on rough horizontal surface (coefficient of friction =0.5),in rest what is speed of block at=20s?
1)15m/s ,2)18m/s,3)25m/s,4)can't be calculated

Anwesha Chakraborty
10 Years agoGrade 12th pass

3 Answers

Vikas TU
10 Years ago
acceleartion = (F-uN)/m =(5t – 0.5 * 100)/10 = (5t – 50)/2 m/s^2
a = dv/dt =(5t – 50)/2 m/s^2
dv = [(5t – 50)/2]dt
integerating both sides,
v = 5t^2/4 – 25t
at t=20s 
put te value nd get d ansr.
 
Anwesha Chakraborty
10 Years ago
Well I tried it out the same way that you have done...and putting the value of t=20 I am getting 0...but answer is given to be 18m/s!
Puja Sahu
8 Years ago
at my end answer is 25m/s so I also think the answer is D – cannot be calculated... I tried every possible way but not getting any answer......................................................